# Disentanglement with Factor Quantized Variational Autoencoders

### v2

@@ A.1 Datasets
Shapes3D: It is a dataset consisting 480,000 images of various 3D geometric objects. It has 6 generative factors: floor hue, wall hue, object hue, object scale, object shape, and camera orientation having 10, 10, 10, 8, 4, and 15 possible values, respectively.
Isaac3D: It is a dataset consisting 737,280 images of a synthetic robot arm holding objects in different configurations. It has 9 generative factors: object shape, robot’s horizontal axis, robot’s vertical axis, camera height, object scale, lighting intensity, lighting direction, object color, and wall color having 3, 8, 5, 4, 4, 4, 6, 4, and 4 possible values, respectively.
MPI3D: It is a dataset consisting 460,800 images of a real robot arm holding objects in different configurations. It has 7 generative factors: object color, object shape, object size, camera height, background color, robot’s horizontal axis, and robot’s vertical axis having 4, 4, 2, 3, 3, 40, and 40 possible values, respectively.
@@ A.2 Hyperparameters
$\mathrm{Conv}_{n}^{(3\times 3)}$ 2D Conv layer (out_ch$=n$, kernel$=3$, stride$=1$, padding$=1$)
$\mathrm{Linear}_{n}$ Linear layer (out_ch$=n$)
$\mathrm{MaxPool}$ 2D Max pooling layer (kernel_size$=2$)
$\mathrm{Upsample}$ 2D upsampling layer (scale_factor$=2$)
$\mathrm{EncResBlock}_{n}$ $3\times(\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(3\times 3)})\to\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(1\times 1)}$ $+$ identity
$\mathrm{DecResBlock}_{n}$ $\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(1\times 1)}\to 3\times(\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(3\times 3)})$ $+$ identity
@@
We run controlled experiments for all models on all datasets, and report the key hyperparameters of the best performing models in Table 3. We use the same coefficient names from the original papers, and try to stick to original hyperparameter values from them. As the complexities of the datasets vary a lot, we naturally obtain the best performances with different hyperparameters.
@@ A.3 Model Description
$n=128$, and $m=128$ for all datasets. For AE, BioAE, QLAE, QVAE, and FactorQVAE, $o=p=d$ where $d$ is the latent dimension. For $\beta$-VAE and FactorVAE, $x$ is encoded into a latent representation with $o=2\times d$ dimensions for the mean and the variance of the posterior distribution, and $p=d$. For dVAE and FactordVAE, $o=p=d*K$ as the encoding part learns the parameters of a distribution over the codebook elements.

### v3

@@ A.1 Datasets
Table 4: Effect of $\beta$ and $\gamma$ on Isaac3D.
$\beta$ $\gamma$ $\mathsf{D}$ $\uparrow$ $\mathsf{I}$ $\uparrow$ $\mathsf{C}$ $\uparrow$ MSE $\downarrow$
$5\!\times\!10^{-5}$ $5\!\times\!10^{-5}$ 0.59 0.75 0.56 $7\!\times\!10^{-5}$
$1\!\times\!10^{-5}$ $5\!\times\!10^{-5}$ 0.56 0.74 0.50 $6\!\times\!10^{-5}$
$5\!\times\!10^{-5}$ $1\!\times\!10^{-5}$ 0.57 0.74 0.52 $6\!\times\!10^{-5}$
$\times\!10^{-5} 1×1$$\times\!10^{-5}$ 0.54 0.73 0.51 $5\!\times\!10^{-5}$
Shapes3D: It is a dataset consisting 480.000 images of various 3D geometric objects. It has 6 generative factors: floor hue, wall hue, object hue, object scale, object shape, and camera orientation having 10, 10, 10, 8, 4, and 15 possible values, respectively.
Isaac3D: It is a dataset consisting 737.280 images of a synthetic robot arm holding objects in different configurations. It has 9 generative factors: object shape, robot’s horizontal axis, robot’s vertical axis, camera height, object scale, lighting intensity, lighting direction, object color, and wall color having 3, 8, 5, 4, 4, 4, 6, 4, and 4 possible values, respectively.
MPI3D: It is a dataset consisting 460.800 images of a real robot arm holding objects in different configurations. It has 7 generative factors: object color, object shape, object size, camera height, background color, robot’s horizontal axis, and robot’s vertical axis having 4, 4, 2, 3, 3, 40, and 40 possible values, respectively.
@@ A.2 Hyperparameters
$\mathrm{Conv}_{n}^{(3\times 3)}$ 2D Conv layer (out_ch=n=n, kernel=3=3, stride=1=1, padding=1=1)
$\mathrm{Linear}_{n}$ Linear layer (out_ch=n=n)
$\mathrm{MaxPool}$ 2D Max pooling layer (kernel_size=2=2)
$\mathrm{Upsample}$ 2D upsampling layer (scale_factor=2=2)
$\mathrm{EncResBlock}_{n}$ $3\times(\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(3\times 3)})\to\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(1\times 1)}$ ++ identity
$\mathrm{DecResBlock}_{n}$ $\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(1\times 1)}\to 3\times(\mathrm{ReLU}\to\mathrm{Conv}_{n}^{(3\times 3)})$ ++ identity
@@
We run controlled experiments for all models on all datasets, and report the key hyperparameters of the best performing models in Table 3. We use the same coefficient names from the original papers, and try to stick to original hyperparameter values from them. As the complexities of the datasets vary a lot, we naturally obtain the best performances with different hypTo examine the effect of $\beta$ and $\gamma$ coefficients in FactorQVAE, we perform a targeted hyperparameter analysis on Isaac3D as in Table 4. As observed in prior disentanglement methods, increasing $\beta$ or $\gamma$ strengthens disentanglement at some cost to reconstruction, while reducing them improves reconstruction quality. For example, with our default setting ($\beta=5\times 10^{-5}$, $\gamma=5\times 10^{-5}$), FactorQVAE achieves $7\times 10^{-5}$ MSE, 0.59 disentanglement, and 0.56 completeness. With smaller values ($\beta=1\times 10^{-5}$, $\gamma=1\times 10^{-5}$), reconstruction improves to $5\times 10^{-5}$ MSE but disentanglement decreases to 0.54 and completeness to 0.51. Across all tested values, FactorQVAE maintains non-trivial disentanglement ($\geq 0.54$), allowing practitioners to dial performance toward reconstruction or factor separation depending on task priorities.
@@ A.3 Model Description
$n=128$, and $m=128$ for all datasets. For AE, BioAE, QLAE, QVAE, and FactorQVAE, $o=p=d$ where $d$ is the latent dimension. For $\beta$-VAE, FactorVAE, and $\alpha$-TCVAE, $x$ is encoded into a latent representation with $o=2\times d$ dimensions for the mean and the variance of the posterior distribution, and $p=d$. For dVAE and FactordVAE, o=p=dKo=p=d*K as the encoding part learns the parameters of a distribution over the codebook elements.